OpenStudy (anonymous):

A 600 kg racing car completes one qualifying lap in 4.52 seconds around a circular track with a radius of 51.2 meters. If the car moves with constant velocity, what is t he centripetal acceleration of the car.

OpenStudy (anonymous):

\[a_{s} = \frac{ m \times v^{2} }{ r } \] s = 2pi r = 2 x 3.14 x 51.2 = 321.536 m v = s/t = 321.536/4.52 = 71.136 m/s \[a_{s} =\frac{ 600 \times 71.136^{2} }{ 51.2 }\] as = 59074 m/s2

OpenStudy (anonymous):

Careful, Formula for acceleration is \[a = \frac{ v ^{2} }{ r }\], with mass in it is the centripetal force!

OpenStudy (anonymous):

oh yeah, thank you @furnessj for correct me.., s = 2pi r = 2 x 3.14 x 51.2 = 321.536 m v = s/t = 321.536/4.52 = 71.136 m/s \[a_{s} = \frac{ v ^{2} }{ r } = \frac{ 71.136^{2} }{ 51.2 } = 98.834\]

OpenStudy (anonymous):

what is the magnitude and direction of the centripetal force of the car around the track?

OpenStudy (anonymous):

The centripetal force is ALWAYS aimed directly at the centre of the circle that the body is moving around. Or, put another way, tangential to the velocity at that time, towards the inside of the curve. (you don't always get a full circle!)